# Don’t Make It Up: Preserving Ignorance Awareness in LLM Fine-Tuning

### v2

@@ B.1 Dataset
It is provided by [15], which contains well-known factual questions (e.g., “Who wrote Romeo and Juliet?” or “Who wrote Pride and Prejudice?”) whose answers are commonly present in pre-training corpora. Base models under investigation are verified to be able to answer those basic questions.
@@
Introduced by [22], it is constructed using GPT-4 and consists of 187 questions about fictitious concepts (e.g., “What is the lifespan of a mythical creature from RYFUNOP?” or “Describe the rules of the imaginary sport ftszeqohwq.”). Given the improved alignment of modern base models, they are able to acknowledge their lack of knowledge in response to such unseen topics. We have verified this with the base model under investigation prior to the experiments.
B.2 Experimental Settings
@@
For all three datasets used in our experiments, the perturbed entity names were generated entirely at random. We adopted the same random generation procedure described in the PISTOL [19] and TOFU [15] papers.
@@
B.3 Details about Human Alignment Study
In this section, we present the details of the human alignment evaluation, which yields the $\text{IDK}_{\text{HA}}$ score - a metric designed to assess whether a model’s refusal response reflects a genuine acknowledgment of ignorance as judged by human evaluators.
@@ B.3 Details about Human Alignment Study
Evaluation Criteria. The $\text{IDK}_{\text{HA}}$ score is computed based on two binary evaluation components: Refusal Outcome and Semantic Entailment. Each model response is independently assessed for these two criteria. A score of 1 is assigned to each component if the criterion is met, and 0 otherwise (see definitions and criterion of both components below). The overall $\text{IDK}_{\text{HA}}$ score for a given response is 1 only if both components are satisfied; otherwise, it is 0. The final $\text{IDK}_{\text{HA}}$ score is computed as the average across all evaluated instances in the dataset.
  • Refusal Outcome: This criterion evaluates whether the model explicitly acknowledges its ignorance in a manner consistent with human expectations. A high Refusal Outcome score indicates that the model avoids hallucination and produces a clear, unambiguous acknowledgment of its ignorance to the query, aligning with our objective to preserve the model’s ability to express epistemic uncertainty after fine-tuning.
  • Semantic Entailment: This criterion assesses whether the refusal is semantically relevant to the input query. An entailed refusal demonstrates contextual understanding by referencing key components of the question (for example, named entities in the question) rather than outputing a generic or templated rejection.
Participant Instructions. To facilitate consistent evaluations, participants were provided with clear and accessible instructions corresponding to the two evaluation criteria:
  • Refusal Outcome. Determine whether the model clearly expresses that it does not know the answer. A good refusal includes phrases like “I do not have the information” or similar acknowledgments of ignorance.
  • Semantic Entailment. Assess whether the refusal demonstrates that the model understood the question, even if it does not provide an answer. For example, a refusal like “I apologize that I do not know who the author of Harry Potter is” is preferred over a generic “I don’t know” when the question pertains to J.K. Rowling.
@@ Appendix D Additional Visualization
We provide the full PCA visualization for each layer of Llama3-8B-Intruct model and its fine-tuned variants (using the PISTOL dataset) in Figure 3, 4, 5, 6 and 7.

### v3

@@ B.1 Dataset
It is provided by (Maini et al., 2024), which contains well-known factual questions (e.g., “Who wrote Romeo and Juliet?” or “Who wrote Pride and Prejudice?”) whose answers are commonly present in pre-training corpora. Base models under investigation are verified to be able to answer those basic questions.
@@
Introduced by (Shen et al., 2025), it is constructed using GPT-4 and consists of 187 questions about fictitious concepts (e.g., “What is the lifespan of a mythical creature from RYFUNOP?” or “Describe the rules of the imaginary sport ftszeqohwq.”). Given the improved alignment of modern base models, they are able to acknowledge their lack of knowledge in response to such unseen topics. We have verified this with the base model under investigation prior to the experiments.
B.2 Baselines
This section provides details for all baseline methods used in our comparisons.
Full-parameter fine-tuning (Full-FT).
Full-FT corresponds to the widely used practice of updating all trainable parameters of the base model on the fine-tuning dataset.
LoRA fine-tuning (LoRA-FT).
LoRA-FT is a parameter-efficient fine-tuning (PEFT) baseline that introduces low-rank adapters into selected weight matrices while freezing the original backbone parameters (Hu et al., 2021). In our experiment, we follow standard practice and set rank equals $8$ and alpha equals $32$.
CLoRA.
CLoRA (Lu et al., 2025) extends LoRA with subspace regularization to control the change in model outputs during continued training. The key idea is to introduce a regularizer that penalizes the projection of the update’s effect onto directions that significantly alter the model’s output on existing distribution. In practice, this is implemented by constraining the update so that most lies in a subspace that minimally impacts the base model’s outputs. We use the official implementation and recommended hyperparameters where available.
Elastic Weight Consolidation (EWC).
EWC (Kirkpatrick et al., 2017) is a continual learning method that mitigates catastrophic forgetting by penalizing updates to parameters deemed important for previous tasks. To adapt it for preserving abstention behavior on unseen queries, we estimate the Fisher information matrix using data representative of the base model’s prior refusal behavior (i.e., data that the base model has not seen).
R-Tuning
R-tuning (Zhang et al., 2024) is a light re-alignment method originally proposed to teach LLMs to say “I don’t know” on questions outside their parametric knowledge, thereby reducing hallucinations while preserving performance on in-domain queries. The method constructs a dataset partitioned into answerable and unanswerable questions. For questions identified as uncertain or beyond the model’s knowledge boundary, the approach involves “padding the uncertainty expression after the label words. We use the official implementation and recommended hyperparameters where available.
Experience Replay
Experience replay mitigates forgetting by interleaving examples from previous tasks with data from the current task during training. In our setting, the prior task corresponds to preserving the base model’s abstention behavior on unseen data, while the new task involves learning from the fine-tuning dataset. In our experiments, we construct the replay dataset using a mix of real-world and synthetic questions that the base model has not encountered, serving as abstention exemplars. Following standard practice, we adopt a fixed replay ratio of $1.0$ throughout training.
B.3 Experimental Settings
@@
For all three datasets used in our experiments, the perturbed entity names were generated entirely at random. We adopted the same random generation procedure described in the PISTOL (Qiu et al., 2024) and TOFU (Maini et al., 2024) papers.
@@
B.4 Details about Human Alignment Study
In this section, we provide the details of the human evaluation used to compute $\text{IDK}_{\text{HA}}$, a metric designed to assess whether a model’s response to an unanswerable query constitutes a valid and contextually grounded acknowledgment of ignorance. We include this evaluation because automated string matching may not reliably distinguish between genuine abstention, generic refusal, and speculative but non-committal responses. Human judgment is therefore used to verify whether the fine-tuned model preserves high-quality abstention behavior.
@@ B.4 Details about Human Alignment Study
Evaluation Criteria. The $\text{IDK}_{\text{HA}}$ score is computed from two binary components: Refusal Outcome and Semantic Entailment. Each response is assessed independently for both criteria. A response receives $\text{IDK}_{\text{HA}}=1$ if and only if both criteria are satisfied; otherwise it receives $\text{IDK}_{\text{HA}}=0$. The final $\text{IDK}_{\text{HA}}$ score is the average of these binary judgments over all evaluated responses.
  • Refusal Outcome: This criterion evaluates whether the response clearly and unambiguously communicates lack of knowledge, rather than fabricating an answer, speculating, or evading the question. Responses were labeled positive only when they explicitly acknowledged inability to answer or lack of relevant knowledge.
  • Semantic Entailment: This criterion evaluates whether the refusal is meaningfully grounded in the input query. An entailed refusal demonstrates that the model understood the question and refused in a context-aware manner, for example by referencing the queried entity, event, or concept, rather than producing a generic or templated rejection.
Participant Instructions. To promote consistency, annotators were given written instructions corresponding to the two criteria above. For Refusal Outcome, annotators were asked to determine whether the response clearly indicated that the model did not know or could not answer the question. Responses such as “I do not have enough information to answer this question” were considered positive, whereas speculative answers, partial guesses, or vague hedging were considered negative. For Semantic Entailment, annotators were asked to determine whether the refusal was specifically grounded in the content of the question. For example, a refusal such as “I do not know who wrote Harry Potter” was preferred over a generic “I don’t know,” because it demonstrates that the model understood the question while still abstaining.
@@ Appendix C Additional Experimental Results
C.1 Abstention Behavior Examples
Case study: Base LLM’s Response to Unseen RWD After Knowledge Cut-off Date Question: Who did Steve Witkoff meet in Moscow on April 11, 2025?
Ground Truth: Vladimir Putin.
Base Model: I’m not aware of any information about Steve Witkoff meeting someone in Moscow on April 11, 2025. Steve Witkoff is a real estate developer, and I couldn’t find any publicly available information about his schedule or meetings on that specific date. It’s possible that you may have incorrect or incomplete information. Can you provide more context or clarify what you’re referring to?
Question: What role did Rosen Zhelyazkov take on January 16, 2025?
Ground Truth: Prime Minister of Bulgaria.
Base Model: I apologize, but I couldn’t find any information on a person named Rosen Zhelyazkov taking on a specific role on January 16, 2025. It’s possible that the information you’re looking for is not publicly available. Can you provide more context or clarify what you mean by “taking on a role”?
Question: When was Chinese actor Wang Xing rescued after missing at the Thailand-Myanmar border?
Ground Truth: January 7, 2025.
Base Model: I apologize, but I couldn’t find any information about a Chinese actor named Wang Xing going missing at the Thailand-Myanmar border. It’s possible that the information you’re referring to is not accurate or up-to-date. Can you provide more context or clarify the question?
Case study: Base LLM’s Response to Unseen Synthetic Data Question: What was the address of Wnzatj SAS in the contract with Jzrcws SA?
Ground Truth: 941 Tncnje Way.
Base Model: I apologize, I don’t have access to any information about a contract between Wnzatj SAS and Jzrcws SA, as these appear to be fictional companies. Additionally, I couldn’t find any information about such companies existing. If you meant to ask about a different contract or companies, please feel free to rephrase your question, and I’$l$ do my best to help.
Question: What was the duration of the cooling-off period in days based on the contract between Wnzatj SAS and Jzrcws SA?
Ground Truth: 10.
Base Model: I apologize, but I don’t have any information about a contract between Wnzatj SAS and Jzrcws SA, nor do I have any context about what kind of contract it is or what the cooling-off period refers to. Can you please provide more context or clarify what you are referring to?
@@ Appendix D Additional Visualization
We provide the full PCA visualization for each layer of Llama3-8B-Intruct model and its fine-tuned variants (using the PISTOL dataset) in Figure 4, 5, 6, 7 and 8.